fox- f 1 idazo 2o if xo g(x) = x² what is tog? Any discontinuity points?
Answer #1

2 > 0 1x)=s 20 g(x) = x² THEIR O x <o -IR Since 2²3,0 AXEIR a) f(22) = 1 THEIR (fog)(x) = flag (x)) = f(x²)=1 their fog(x) =plz do like